# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

### v1

@@ 4. Multilingual Hate Speech Datasets
(Paul et al., 2023) |
2023 |
cyberbullying |
en, hi |
22k |
1 |
¡50 |
(Jeong et al., 2022) |
2022 |
offense |
NAVER-news, Youtube |
ko |
40,429 |
1,2,3 |
¡10 |
(Jiang et al., 2022) |
2022 |
Sexism |
zh |
8,969 |
1,2,3 |
¡50 |
(Salaam et al., 2022) |
2022 |
offense |
en, fr, de, es |
37,221 |
1 |
¡10 |
(Deng et al., 2022) |
2022 |
offense |
Weibo, Zhihu |
zh |
37,480 |
1,2 |
¡50 |
(Gupta et al., 2022) |
2022 |
abuse |
ShareChat |
hi, kn, ml, ta, te |
92,881 |
1 |
¡10 |
(Chakravarthi et al., 2022) |
2022 |
offense |
Youtube |
ta |
60k |
1,3 |
¡50 |
(Arango Monnar et al., 2022) |
2022 |
offense |
es |
9,834 |
2 |
¡10 |
(Bhattacharya et al., 2020) |
2022 |
aggressiveness, misogyny |
Youtube |
bn |
25k |
1 |
¡100 |
(Kumar and Sachdeva, 2022) |
2022 |
cyberbullying |
Facebook, Twitter |
en, hi |
6,500 |
1 |
¡50 |
(Pavlopoulos et al., 2021) |
2021 |
toxicity |
Civil Comments |
en |
10,629 |
1,2 |
¡100 |
(Burtenshaw and Kestemont, 2021) |
2021 |
toxicity |
Ask.fm |
nl |
10,189 |
2 |
¡10 |
(Hande et al., 2021b) |
2021 |
offense |
Youtube |
kn, ml, ta |
71,691 |
1,3 |
¡20 |
(Mathew et al., 2021) |
2021 |
hate speech |
Gab, Twitter |
en |
20,148 |
1,2 |
¡500 |
(Romim et al., 2021) |
2021 |
hate speech |
Facebook, Youtube |
bn |
30k |
1 |
¡50 |
(Modha et al., 2021) |
2021 |
hate speech, offense |
en, hi, mr |
13,755 |
1,4 |
¡100 |
(Amjad et al., 2022) |
2021 |
abuse |
ur |
8,400 |
1 |
¡50 |
(Vidgen et al., 2020) |
2021 |
hate speech |
open-source platform |
en |
41,255 |
1,2 |
¡100 |
(Gaikwad et al., 2021) |
2021 |
offense |
mr |
2499 |
1,2,3 |
¡50 |
(Khan et al., 2021) |
2021 |
hate speech, offense |
roman ur |
5k |
1 |
¡50 |
(Mubarak et al., 2021) |
2021 |
offense |
ar |
10k |
1,2 |
¡250 |
(Pitenis et al., 2020) |
2020 |
offense |
el |
4,779 |
1 |
¡250 |
(Charitidis et al., 2020) |
2020 |
hate speech |
en, fr, de, el, es |
264,035 |
1 |
¡50 |
(Glavaš et al., 2020) |
2020 |
hate speech, abuse |
Facebook, Fox news, Twitter, Wikipedia |
sq, hr, en, de, ru, tr |
109,955 |
1 |
¡50 |
(Shekhar et al., 2020) |
2020 |
offense |
24sata, Eesti Ekspress, Veˇcernji List |
hr, et |
62.6m |
¡10 |
(Akhter et al., 2020) |
2020 |
offense, abuse |
Youtube |
Roman ur, ur |
12,171 |
1 |
¡100 |
(Rizwan et al., 2020) |
2020 |
hate speech, offense |
Roman ur |
10,012 |
1,2 |
¡50 |
(Mandl et al., 2021) |
2020 |
offense |
Twitter, Youtube |
en, de, hi, ml, ta |
15,047 |
1,2 |
¡250 |
(Sanguinetti et al., 2020) |
2020 |
hate speech |
it |
12,081 |
1 |
¡100 |
(Kennedy et al., 2020) |
2020 |
hate speech |
Gab |
en |
27,655 |
1 |
¡250 |
(Çöltekin, 2020) |
2020 |
offense |
tr |
36,232 |
1,3 |
¡250 |
(Kumar et al., 2020) |
2020 |
aggressiveness |
Youtube |
bn, en, hi |
15k |
1 |
¡250 |
(Sigurbergsson and Derczynski, 2020) |
2020 |
abuse, offense |
Facebook, Reddit, Twitter |
da |
3,600 |
1,3 |
¡250 |
(Hande et al., 2020) |
2020 |
offense |
Youtube |
kn |
7,671 |
1,3 |
¡100 |
(Zampieri et al., 2020) |
2020 |
offense |
ar, da, en, el, tr |
9m |
1,3 |
¡500 |
(Ljubešić et al., 2019) |
2019 |
offense |
en, sl |
22,877 |
1,2 |
¡50 |
(Arata, 2019) |
2019 |
cyberbullying |
ja |
4,096 |
1 |
¡10 |
(Arango et al., 2019) |
2019 |
hate speech |
en |
14,949 |
1 |
¡250 |
(Ibrohim and Budi, 2019b) |
2019 |
abuse, hate speech |
id |
5,561 |
1,2,3 |
¡250 |
(Mulki et al., 2019) |
2019 |
abuse, hate speech, toxicity |
at |
5,846 |
1 |
¡250 |
(Pereira-Kohatsu et al., 2019) |
2019 |
hate speech |
es |
6k |
1 |
¡100 |
(Fortuna et al., 2019) |
2019 |
hate speech, offense |
pt |
5,668 |
1 |
¡100 |
(Poletto et al., 2021) |
2019 |
cyberbullying, toxicity |
pl |
11,041 |
1,2 |
¡50 |
(Strus̈ et al., 2019) |
2019 |
offense |
de |
7,025 |
1,2 |
¡250 |
(Basile et al., 2019) |
2019 |
aggressiveness |
en |
19,600 |
1,3 |
¡750 |
(Zampieri et al., 2019) |
2019 |
offense |
en, es |
14k |
1,3 |
¡700 |
(Mandl et al., 2019) |
2019 |
hate speech, offense |
Twitter, Facebook |
en, de, hi |
17,657 |
1,2,3 |
¡100 |
(Chung et al., 2019) |
2019 |
hate speech |
en, fr, it |
4,078 |
1,2 |
¡250 |
(Ousidhoum et al., 2019) |
2019 |
hate speech |
ar, en, fr |
13,014 |
1,2,3 |
¡250 |
(Chiril et al., 2019) |
2019 |
sexism, misogyny |
fr |
3,085 |
1 |
¡50 |
(Bosco et al., 2018) |
2018 |
hate speech |
Facebook, Twitter |
en, it |
4k |
1 |
¡250 |
(de Gibert et al., 2018) |
2018 |
hate speech |
Stormfront |
en |
10,568 |
1 |
¡500 |
(Ptaszyński et al., 2018) |
2018 |
cyberbullying |
Formspring.me |
en |
12,772 |
1 |
¡50 |
(Alakrot et al., 2018) |
2018 |
abuse, offense |
Youtube |
ar |
167,549 |
1 |
¡100 |
(Bohra et al., 2018) |
2018 |
hate speech |
en, hi |
4,575 |
1 |
¡250 |
(ElSherief et al., 2018) |
2018 |
hate speech |
en |
27,330 |
1,2 |
¡250 |
(Fersini et al., 2018a) |
2018 |
misogyny |
en, it |
20k |
1,2,3 |
¡250 |
(Mathur et al., 2018a) |
2018 |
abuse, hate speech, offense |
en, hi |
17,698 |
1,2 |
¡100 |
(Kumar et al., 2018) |
2018 |
aggressiveness |
Facebook, Twitter |
en, hi |
39k |
1,2,3 |
¡250 |
(Mathur et al., 2018b) |
2018 |
offense |
en, hi |
3,679 |
1,2 |
¡250 |
(Sanguinetti et al., 2018) |
2018 |
aggressiveness, hate speech, offense |
it |
6k |
2,4 |
¡250 |
(Founta et al., 2018) |
2018 |
abuse, hate speech |
en |
80k |
1,2 |
¡550 |
(Wiegand et al., 2018) |
2018 |
hate speech, offense |
de |
8,541 |
1,2 |
¡500 |
(Fersini et al., 2018b) |
2018 |
misogyny |
en, es |
8,115 |
1,2,3 |
¡250 |
(Golbeck et al., 2017) |
2017 |
harassment, offense, racism |
BlockTogether, Twitter |
en |
35k |
1 |
¡100 |
(Devyatkin et al., 2017) |
2017 |
offense |
ru |
493 |
1,2 |
¡50 |
(Gao and Huang, 2017) |
2017 |
hate speech |
Fox news |
en |
1,528 |
1 |
¡250 |
(Wulczyn et al., 2017) |
2017 |
hate speech |
Wikipedia |
en |
115,737 |
1 |
¡750 |
(de Pelle and Moreira, 2017) |
2017 |
hate speech, offense |
g1.globo.com |
pt |
1,250 |
1 |
¡100 |
(Ross et al., 2017) |
2017 |
hate speech, offense |
de |
541 |
1,2,4 |
¡500 |
(Alfina et al., 2017) |
2017 |
hate speech |
id |
713 |
1 |
¡250 |
(Fortuna, 2017) |
2017 |
hate speech |
pt |
5,668 |
1,2 |
¡50 |
(Bretschneider and Peters, 2017) |
2017 |
hate speech, offense |
de |
5,836 |
1,3 |
¡100 |
(Davidson et al., 2017) |
2017 |
offense |
en |
24,802 |
1,2 |
¡2200 |
(Mubarak et al., 2017) |
2017 |
abuse, hate speech, offense |
ar |
32k |
1 |
¡500 |
(Waseem and Hovy, 2016) |
2016 |
hate speech |
en |
16,914 |
1 |
¡1550 |
(Van Hee et al., 2015) |
2015 |
cyberbullying |
Ask.fm |
nl |
85,462 |
2,4 |
¡250 |
(Dybala et al., 2010) |
2010 |
cyberbullying |
informal websites of Japanese secondary schools |
ja |
2,999 |
1 |
¡100 |
Topics
The topic distribution among datasets shows that “offensiveness” and “hate speech” are the two most frequently addressed topics, accounting for 34.23% and 32.43% respectively. Some cover general offense, such as abusiveness, toxicity, cyberbullying, harassment and aggressiveness. In addition, some datasets narrow their scope to targeted issues, like sexism, misogyny and racism.
Data Sources
Most studies tend to collect data from openly accessible social media sources such as Twitter and Facebook. Our review of data sources reveals a pronounced reliance on popular social media platforms, with Twitter notably encompassing 47.19% of the datasets. This is followed by Youtube and Facebook, which amount to 11.24% and 10.11% respectively. While these mainstream platforms dominate, news websites like Fox News and NAVER news, and open forums like Reddit and 2ch, also contribute to the multilingual datasets. And there is a rich diversity in the data source origins. Sources like Weibo (Chinese), 2ch (Russian), g1.globo (Brazilian Portuguese) and Eesti Ekspress (Estonian) represent some non-English speaking regions as more culturally specific platforms. Some datasets originate from platforms known for controversial content or specific user bases, such as Stormfront and alt-right websites.
Languages and Families
Datasets cover 32 languages across 10 language families, where one dataset may contain more than one language or family. We show the distribution of languages and their language families in Figure 3. The distribution of languages in the datasets highlights a representative language of English, accounting for 25.95% of the data, followed by Hindi at 9.16%, and German at 6.87%. Notably, the datasets exhibit a relatively balanced representation across several languages, such as Spanish, Arabic, Italian, French, Portuguese and Turkish, each ranging from around 2% to 5%. This distribution reveals a strong emphasis on Indo-European languages in hate speech studies, complemented by a notable presence of Afro-Asiatic languages (represented primarily by Arabic), with other language families being less represented and restricting cross-lingual research. Some datasets (14.6%) are also dedicated to code-mixed content.
Dataset Size
Over half of the datasets, accounting for 51.20%, contain - instances. This is closely followed by 35.40% of datasets that have instances between and . On the other hand, large datasets are rare, with only one dataset’s size exceeding , and two between and . We can observe that the majority of datasets fall within the smaller size ranges, which indicates the challenges in collecting and manually annotating large-scale labelled datasets, as well as model generalisability for cross-lingual hate speech research.
Data Labelling and Distributions
According to diverse annotation schemes, four different types of labels are often employed: (i) binary labels, (ii) fine-grained categories of offensive content, (iii) attack targets, and (iv) intensity scores. A majority of datasets, 76 in total, utilise the simpler binary labelling, of which 36 solely provide binary labels. There are also datasets combining binary labels with either fine-grained categories (20), attack targets (9), or intensity scores (1). A smaller number of datasets (10) combine binary labels, fine-grained categories, and attack targets. Interestingly, only four datasets consider intensity scores, suggesting that quantifying the severity of offensive content is less common in current research.
Dataset Availability
We find 66 out of 82 datasets (80%) are publicly accessible, with the other 16 unavailable, in some cases with the possibility of contacting the authors to request the data.
Competitions and Shared Tasks
The rising concern surrounding online abuse has led to the establishment of numerous competitions and shared tasks within both national (e.g., GermEval, Evalita, IberLEF) and international (e.g., SemEval) evaluation campaigns. These open scientific competitions release benchmark datasets and invite participants to submit detection results and detailed system reports (Poletto et al., 2021). Table 3 lists 24 such competitions and shared tasks from 2018 to 2023, each addressing different facets of offensive content across multiple languages.
Prominent topics include offense and hate speech. Specifically, GermEval competitions in 2018 (Wiegand et al., 2018) and 2019 (Strus̈ et al., 2019) centre around identifying offensive content in German tweets. OffensEval, initially in 2019 focused only on English content (Zampieri et al., 2019), expanded to include multiple languages in the subsequent edition in 2020 (Zampieri et al., 2020). The task in OSACT4 (Mubarak et al., 2020), the 4th Workshop on Open-Source Arabic Corpora and Processing Tools, delves into the challenges of offensive language detection in Arabic, especially its dialectal forms. HaSpeeDe has consistently addressed hate speech, with its iterations focusing on various general and particular hate speech topics (Bosco et al., 2018; Sanguinetti et al., 2020; Lai et al., 2023). For instance, HaSpeeDe (Bosco et al., 2018) and HaSpeeDe 2 (Sanguinetti et al., 2020) concentrate on hate speech against immigrants and Muslims, while HaSpeeDe 3 (Lai et al., 2023) explores hate speech in strongly polarised debates, in particular concerning political and religious topics. Similarly, the HASOC series, recurring annually from 2019 to 2023, delves into hate speech and offensive content identification across various languages (Mandl et al., 2019, 2021; Modha et al., 2021; Satapara et al., 2023). TRAC workshops, held in 2018 (Kumar et al., 2018) and 2020 (Kumar et al., 2020), tackled trolling, aggression, and cyberbullying. Other competitions, like PolEval (Ptaszynski et al., 2019) and UrduThreat (Amjad et al., 2022), focus on specific languages (such as Polish and Urdu) or topics (e.g., cyberbullying and threatening languages). The Jigsaw Multilingual Toxic Comment Classification competition on Kaggle attracted over 1600 participants to build multilingual models using English-only training data. Another Kaggle competition, IIIT-D Multilingual Abusive Comment Identification, focuses on abusive comments in Indic languages. Additionally, some tasks focus on more specific topics, such as sexism and misogyny. AMI is organised by IberEval (Fersini et al., 2018b) and Evalita (Fersini et al., 2018a) in 2018, specifically addressing Spanish and Italian misogynous content respectively. EXIST is held for three consecutive years from 2021 to 2023 and aims to identify sexism in social networks (Rodríguez-Sánchez et al., 2021, 2022; Plaza et al., 2023).
The majority of these benchmark datasets originates from social media platforms like Twitter, and often conduct annotations that go beyond simple binary labels to finer-grained ones, enabling to further and broaden research in cross-lingual hate speech detection.
@@
Identification of Offensive Language |
2018 |
offense |
de |
20 | |
AMI@IberEval (Fersini et al., 2018b) |
Automatic Misogyny Identification |
2018 |
misogyny |
en, es |
11 |
AMI@Evalita (Fersini et al., 2018a) |
Automatic Misogyny Identification |
2018 |
misogyny |
en, it |
16 |
HaSpeeDe@Evalita (Bosco et al., 2018) |
Hate Speech Detection |
2018 |
hate speech |
it |
9 |
@@
Aggression Identification of Hindi-English Code-mixed Data |
2018 |
aggressiveness |
en, hi |
30 | |
HatEval@SemEval (Basile et al., 2019) |
@@
Multilingual Detection of Hate Speech Against Immigrants and Women in Twitter |
2019 |
hate speech |
en, es |
74 |
HASOC@FIRE (Mandl et al., 2019) |
@@
Offensive Content Identification in Indo-European Languages |
2019 |
hate speech, offense |
en, hi, de |
37 |
Task2@GermEval (Strus̈ et al., 2019) |
Identification of Offensive Language |
2019 |
offense |
de |
13 |
Task6@PolEval (Ptaszynski et al., 2019) |
@@
Automatic Cyberbullying Detection in Polish Twitter |
2019 |
cyberbullying |
pl |
9 |
OffensEval@SemEval (Zampieri et al., 2020) |
@@
Multilingual Offensive Language Identification in Social Media |
2020 |
offense |
en, ar, da, el, tr |
145 |
HASOC@FIRE (Mandl et al., 2021) |
@@
Offensive Language Identification in Tamil, Malayalam, Hindi, English and German |
2020 |
hate speech, offense |
en, hi, de, ta, ml |
40+ |
HaSpeeDe@Evalita (Sanguinetti et al., 2020) |
Hate Speech Detection |
2020 |
hate speech |
it |
14 |
Aggression and Gendered Aggression Identification |
2020 |
aggressiveness |
en, hi, bn |
19 | |
Jigsaw Multilingual Toxic Comment Classification |
2020 |
toxicity |
en, es, tr, pt |
1621 | |
Arabic Offensive Language Detection |
2020 |
offense |
ar |
27 | |
HASOC@FIRE (Modha et al., 2021) |
@@
Offensive Content Identification in English and Indo-Aryan Languages and Conversational Hate Speech |
2021 |
hate speech, offense |
en, hi, mr |
65 |

### v2

@@ 4. Multilingual Hate Speech Datasets
| (Paul et al., 2023) | 2023 | cyberbullying | en, hi | 22k | 1 | ¡50 | ✓ | ✗ | |
| (Jeong et al., 2022) | 2022 | offense | NAVER-news, Youtube | ko | 40,429 | 1,2,3 | ¡10 | ✗ | ✓ |
| (Jiang et al., 2022) | 2022 | Sexism | zh | 8,969 | 1,2,3 | ¡50 | ✗ | ✓ | |
| (Salaam et al., 2022) | 2022 | offense | - | en, fr, de, es | 37,221 | 1 | ¡10 | ✓ | ✓ |
| (Deng et al., 2022) | 2022 | offense | Weibo, Zhihu | zh | 37,480 | 1,2 | ¡50 | ✗ | ✓ |
| (Gupta et al., 2022) | 2022 | abuse | ShareChat | hi, kn, ml, ta, te | 92,881 | 1 | ¡10 | ✓ | ✓ |
| (Chakravarthi et al., 2022) | 2022 | offense | Youtube | ta | 60k | 1,3 | ¡50 | ✓ | ✓ |
| (Arango Monnar et al., 2022) | 2022 | offense | es | 9,834 | 2 | ¡10 | ✗ | ✓ | |
| (Bhattacharya et al., 2020) | 2022 | aggressiveness, misogyny | Youtube | bn | 25k | 1 | ¡100 | ✗ | ✓ |
| (Kumar and Sachdeva, 2022) | 2022 | cyberbullying | Facebook, Twitter | en, hi | 6,500 | 1 | ¡50 | ✓ | ✗ |
| (Pavlopoulos et al., 2021) | 2021 | toxicity | Civil Comments | en | 10,629 | 1,2 | ¡100 | ✗ | ✓ |
| (Burtenshaw and Kestemont, 2021) | 2021 | toxicity | Ask.fm | nl | 10,189 | 2 | ¡10 | ✗ | ✗ |
| (Hande et al., 2021b) | 2021 | offense | Youtube | kn, ml, ta | 71,691 | 1,3 | ¡20 | ✓ | ✓ |
| (Mathew et al., 2021) | 2021 | hate speech | Gab, Twitter | en | 20,148 | 1,2 | ¡500 | ✗ | ✓ |
| (Romim et al., 2021) | 2021 | hate speech | Facebook, Youtube | bn | 30k | 1 | ¡50 | ✗ | ✗ |
| (Modha et al., 2021) | 2021 | hate speech, offense | en, hi, mr | 13,755 | 1,4 | ¡100 | ✗ | ✓ | |
| (Amjad et al., 2022) | 2021 | abuse | ur | 8,400 | 1 | ¡50 | ✗ | ✓ | |
| (Vidgen et al., 2020) | 2021 | hate speech | open-source platform | en | 41,255 | 1,2 | ¡100 | ✗ | ✓ |
| (Gaikwad et al., 2021) | 2021 | offense | mr | 2499 | 1,2,3 | ¡50 | ✗ | ✓ | |
| (Khan et al., 2021) | 2021 | hate speech, offense | roman ur | 5k | 1 | ¡50 | ✗ | ✗ | |
| (Mubarak et al., 2021) | 2021 | offense | ar | 10k | 1,2 | ¡250 | ✗ | ✓ | |
| (Pitenis et al., 2020) | 2020 | offense | el | 4,779 | 1 | ¡250 | ✗ | ✓ | |
| (Charitidis et al., 2020) | 2020 | hate speech | en, fr, de, el, es | 264,035 | 1 | ¡50 | ✗ | ✓ | |
| (Glavaš et al., 2020) | 2020 | hate speech, abuse | Facebook, Fox news, Twitter, Wikipedia | sq, hr, en, de, ru, tr | 109,955 | 1 | ¡50 | ✗ | ✓ |
| (Shekhar et al., 2020) | 2020 | offense | 24sata, Eesti Ekspress, Veˇcernji List | hr, et | 62.6m | - | ¡10 | ✗ | ✗ |
| (Akhter et al., 2020) | 2020 | offense, abuse | Youtube | Roman ur, ur | 12,171 | 1 | ¡100 | ✗ | ✓ |
| (Rizwan et al., 2020) | 2020 | hate speech, offense | Roman ur | 10,012 | 1,2 | ¡50 | ✗ | ✓ | |
| (Mandl et al., 2021) | 2020 | offense | Twitter, Youtube | en, de, hi, ml, ta | 15,047 | 1,2 | ¡250 | ✓ | ✓ |
| (Sanguinetti et al., 2020) | 2020 | hate speech | it | 12,081 | 1 | ¡100 | ✗ | ✓ | |
| (Kennedy et al., 2020) | 2020 | hate speech | Gab | en | 27,655 | 1 | ¡250 | ✗ | ✓ |
| (Çöltekin, 2020) | 2020 | offense | tr | 36,232 | 1,3 | ¡250 | ✗ | ✓ | |
| (Kumar et al., 2020) | 2020 | aggressiveness | Youtube | bn, en, hi | 15k | 1 | ¡250 | ✗ | ✓ |
| (Sigurbergsson and Derczynski, 2020) | 2020 | abuse, offense | Facebook, Reddit, Twitter | da | 3,600 | 1,3 | ¡250 | ✗ | ✓ |
| (Hande et al., 2020) | 2020 | offense | Youtube | kn | 7,671 | 1,3 | ¡100 | ✓ | ✗ |
| (Zampieri et al., 2020) | 2020 | offense | ar, da, en, el, tr | 9m | 1,3 | ¡500 | ✗ | ✓ | |
| (Ljubešić et al., 2019) | 2019 | offense | en, sl | 22,877 | 1,2 | ¡50 | ✓ | ✗ | |
| (Arata, 2019) | 2019 | cyberbullying | ja | 4,096 | 1 | ¡10 | ✗ | ✗ | |
| (Arango et al., 2019) | 2019 | hate speech | en | 14,949 | 1 | ¡250 | ✗ | ✓ | |
| (Ibrohim and Budi, 2019b) | 2019 | abuse, hate speech | id | 5,561 | 1,2,3 | ¡250 | ✗ | ✓ | |
| (Mulki et al., 2019) | 2019 | abuse, hate speech, toxicity | at | 5,846 | 1 | ¡250 | ✗ | ✗ | |
| (Pereira-Kohatsu et al., 2019) | 2019 | hate speech | es | 6k | 1 | ¡100 | ✗ | ✓ | |
| (Fortuna et al., 2019) | 2019 | hate speech, offense | pt | 5,668 | 1 | ¡100 | ✗ | ✓ | |
| (Poletto et al., 2021) | 2019 | cyberbullying, toxicity | pl | 11,041 | 1,2 | ¡50 | ✗ | ✓ | |
| (Strus̈ et al., 2019) | 2019 | offense | de | 7,025 | 1,2 | ¡250 | ✗ | ✓ | |
| (Basile et al., 2019) | 2019 | aggressiveness | en | 19,600 | 1,3 | ¡750 | ✗ | ✓ | |
| (Zampieri et al., 2019) | 2019 | offense | en, es | 14k | 1,3 | ¡700 | ✗ | ✓ | |
| (Mandl et al., 2019) | 2019 | hate speech, offense | Twitter, Facebook | en, de, hi | 17,657 | 1,2,3 | ¡100 | ✗ | ✓ |
| (Chung et al., 2019) | 2019 | hate speech | en, fr, it | 4,078 | 1,2 | ¡250 | ✗ | ✓ | |
| (Ousidhoum et al., 2019) | 2019 | hate speech | ar, en, fr | 13,014 | 1,2,3 | ¡250 | ✗ | ✓ | |
| (Chiril et al., 2019) | 2019 | sexism, misogyny | fr | 3,085 | 1 | ¡50 | ✗ | ✗ | |
| (Bosco et al., [n. d.]) | 2018 | hate speech | Facebook, Twitter | en, it | 4k | 1 | ¡250 | ✗ | ✓ |
| (de Gibert et al., 2018) | 2018 | hate speech | Stormfront | en | 10,568 | 1 | ¡500 | ✗ | ✓ |
| (Ptaszyński et al., 2018) | 2018 | cyberbullying | Formspring.me | en | 12,772 | 1 | ¡50 | ✗ | ✗ |
| (Alakrot et al., 2018) | 2018 | abuse, offense | Youtube | ar | 167,549 | 1 | ¡100 | ✗ | ✓ |
| (Bohra et al., 2018) | 2018 | hate speech | en, hi | 4,575 | 1 | ¡250 | ✓ | ✓ | |
| (ElSherief et al., 2018) | 2018 | hate speech | en | 27,330 | 1,2 | ¡250 | ✗ | ✓ | |
| (Fersini et al., 2018a) | 2018 | misogyny | en, it | 20k | 1,2,3 | ¡250 | ✗ | ✓ | |
| (Mathur et al., 2018a) | 2018 | abuse, hate speech, offense | en, hi | 17,698 | 1,2 | ¡100 | ✗ | ✓ | |
| (Kumar et al., 2018) | 2018 | aggressiveness | Facebook, Twitter | en, hi | 39k | 1,2,3 | ¡250 | ✓ | ✓ |
| (Mathur et al., 2018b) | 2018 | offense | en, hi | 3,679 | 1,2 | ¡250 | ✓ | ✗ | |
| (Sanguinetti et al., 2018) | 2018 | aggressiveness, hate speech, offense | it | 6k | 2,4 | ¡250 | ✗ | ✓ | |
| (Founta et al., 2018) | 2018 | abuse, hate speech | en | 80k | 1,2 | ¡550 | ✗ | ✓ | |
| (Wiegand et al., 2018) | 2018 | hate speech, offense | de | 8,541 | 1,2 | ¡500 | ✗ | ✓ | |
| (Fersini et al., 2018b) | 2018 | misogyny | en, es | 8,115 | 1,2,3 | ¡250 | ✗ | ✓ | |
| (Golbeck et al., 2017) | 2017 | harassment, offense, racism | BlockTogether, Twitter | en | 35k | 1 | ¡100 | ✗ | ✓ |
| (Devyatkin et al., 2017) | 2017 | offense | ru | 493 | 1,2 | ¡50 | ✗ | ✗ | |
| (Gao and Huang, 2017) | 2017 | hate speech | Fox news | en | 1,528 | 1 | ¡250 | ✗ | ✓ |
| (Wulczyn et al., 2017) | 2017 | hate speech | Wikipedia | en | 115,737 | 1 | ¡750 | ✗ | ✓ |
| (de Pelle and Moreira, 2017) | 2017 | hate speech, offense | g1.globo.com | pt | 1,250 | 1 | ¡100 | ✗ | ✓ |
| (Ross et al., 2017) | 2017 | hate speech, offense | de | 541 | 1,2,4 | ¡500 | ✗ | ✓ | |
| (Alfina et al., 2017) | 2017 | hate speech | id | 713 | 1 | ¡250 | ✗ | ✓ | |
| (Fortuna, 2017) | 2017 | hate speech | pt | 5,668 | 1,2 | ¡50 | ✗ | ✓ | |
| (Bretschneider and Peters, 2017) | 2017 | hate speech, offense | de | 5,836 | 1,3 | ¡100 | ✗ | ✓ | |
| (Davidson et al., 2017) | 2017 | offense | en | 24,802 | 1,2 | ¡2200 | ✗ | ✓ | |
| (Mubarak et al., 2017) | 2017 | abuse, hate speech, offense | ar | 32k | 1 | ¡500 | ✗ | ✓ | |
| (Waseem and Hovy, 2016) | 2016 | hate speech | en | 16,914 | 1 | ¡1550 | ✗ | ✓ | |
| (Van Hee et al., 2015) | 2015 | cyberbullying | Ask.fm | nl | 85,462 | 2,4 | ¡250 | ✗ | ✗ |
| (Dybala et al., 2010) | 2010 | cyberbullying | informal websites of Japanese secondary schools | ja | 2,999 | 1 | ¡100 | ✗ | ✗ |
Topics. The topic distribution among datasets shows that “offensiveness” and “hate speech” are the two most frequently addressed topics, accounting for 34.23% and 32.43% respectively. Some cover general offense, such as abusiveness, toxicity, cyberbullying, harassment and aggressiveness. In addition, some datasets narrow their scope to targeted issues, like sexism, misogyny and racism.
Data Sources. Most studies collect data from accessible social media such as Twitter and Facebook. Our review reveals a pronounced reliance on popular social media. Twitter encompasses 47.19% of the datasets, followed by Youtube and Facebook (11.24% and 10.11% respectively). While these mainstream platforms dominate, news websites like Fox News and NAVER news, and open forums like Reddit and 2ch, also contribute to the multilingual datasets. And there is a rich diversity in the data source origins. Sources like Weibo (Chinese), 2ch (Russian), g1.globo (Brazilian Portuguese) and Eesti Ekspress (Estonian) represent non-English speaking regions as more culturally specific platforms. Some datasets originate from platforms known for controversial content or specific user bases, such as Stormfront and alt-right websites.
Languages and Families. Datasets cover 32 languages across 10 language families, where one dataset may contain more than one language or family. We show the distribution of languages and families in Figure 3. It highlights a representative language of English, accounting for 25.95% of the data, followed by Hindi at 9.16%, and German at 6.87%. Notably, it exhibits a relatively balanced representation across several languages, such as Spanish, Arabic, Italian, French, Portuguese and Turkish, each ranging from around 2% to 5%. It also reveals a strong emphasis on Indo-European languages in HS studies, complemented by a notable presence of Afro-Asiatic languages (represented primarily by Arabic), with other language families being less represented. Some datasets (14.6%) are also dedicated to code-mixed content.
Dataset Size. Over half of the datasets, accounting for 51.20%, contain - instances, closely followed by 35.40% that have instances between and . Besides, large datasets are rare, with only one exceeding , and two between and . The majority of datasets fall within the smaller size ranges, which indicates the challenges in collecting and manually annotating large-scale datasets, as well as model generalisability for cross-lingual HS.
Data Labelling and Distributions. According to diverse annotation schemes, four different types of labels are often employed: (i) binary labels, (ii) fine-grained categories of offensive content, (iii) attack targets, and (iv) intensity scores. A majority of datasets, 76 in total, utilise the simpler binary labelling, of which 36 solely provide binary labels. There are also datasets combining binary labels with either fine-grained categories (20), attack targets (9), or intensity scores (1). A smaller number of datasets (10) combine binary labels, fine-grained categories, and attack targets. Interestingly, only four datasets consider intensity scores, suggesting that quantifying the severity of offensive content is less common in current research.
Dataset Availability. We find 66 out of 82 datasets (80%) are publicly accessible, with the other 16 unavailable, in some cases with the possibility of contacting the authors to request the data.
@@
| GermEval (Wiegand et al., 2018) | Identification of Offensive Language | 2018 | offense | de | 20 |
| AMI@IberEval (Fersini et al., 2018b) | Automatic Misogyny Identification | 2018 | misogyny | en, es | 11 |
| AMI@Evalita (Fersini et al., 2018a) | Automatic Misogyny Identification | 2018 | misogyny | en, it | 16 |
| HaSpeeDe@Evalita (Bosco et al., [n. d.]) | Hate Speech Detection | 2018 | hate speech | it | 9 |
@@
| TRAC-1 (Kumar et al., 2018) | Aggression Identification of Hindi-English Code-mixed Data | 2018 | aggressiveness | en, hi | 30 |
@@
| HatEval@SemEval (Basile et al., 2019) | Multilingual Detection of Hate Speech Against Immigrants and Women in Twitter | 2019 | hate speech | en, es | 74 |
@@
| HASOC@FIRE (Mandl et al., 2019) | Offensive Content Identification in Indo-European Languages | 2019 | hate speech, offense | en, hi, de | 37 |
| Task2@GermEval (Strus̈ et al., 2019) | Identification of Offensive Language | 2019 | offense | de | 13 |
@@
| Task6@PolEval (Ptaszynski et al., 2019) | Automatic Cyberbullying Detection in Polish Twitter | 2019 | cyberbullying | pl | 9 |
@@
| OffensEval@SemEval (Zampieri et al., 2020) | Multilingual Offensive Language Identification in Social Media | 2020 | offense | en, ar, da, el, tr | 145 |
@@
| HASOC@FIRE (Mandl et al., 2021) | Offensive Language Identification in Tamil, Malayalam, Hindi, English and German | 2020 | hate speech, offense | en, hi, de, ta, ml | 40+ |
| HaSpeeDe@Evalita (Sanguinetti et al., 2020) | Hate Speech Detection | 2020 | hate speech | it | 14 |
| TRAC-2 (Kumar et al., 2020) | Aggression and Gendered Aggression Identification | 2020 | aggressiveness | en, hi, bn | 19 |
| Jigsaw Toxic@Kaggle | Jigsaw Multilingual Toxic Comment Classification | 2020 | toxicity | en, es, tr, pt | 1621 |
| OSACT4 (Mubarak et al., 2020) | Arabic Offensive Language Detection | 2020 | offense | ar | 27 |
@@
| HASOC@FIRE (Modha et al., 2021) | Offensive Content Identification in English and Indo-Aryan Languages and Conversational Hate Speech | 2021 | hate speech, offense | en, hi, mr | 65 |